OpenStudy (anonymous):

A recipe that makes 6 servings calls for 1/3 cup of sugar. Arlene modifies the recipe so that it can serve 9. How many cups of sugar does Arlene need?

OpenStudy (kittiwitti1):

Let's make this info into ratios. 6 servings: 1/3 cup sugar 9 servings: ? cup(s) sugar Now we turn this into fraction-form ratios:\[\frac{6\text{ servings}}{\frac{1}{3}\text{ cup sugar}}=\frac{9\text{ servings}}{\text{? cups of sugar}}\]Can you solve from here?

OpenStudy (kittiwitti1):

Hint: cross-multiply

OpenStudy (kittiwitti1):

Are you still stuck?

OpenStudy (anonymous):

3?

OpenStudy (kittiwitti1):

You solved it? :o

OpenStudy (anonymous):

I did? lol awesome

OpenStudy (kittiwitti1):

Well, I don't know if your answer is right. I will check, hold on :)

OpenStudy (coconutjj):

I don't think its correct... 6x3 = 18, 9/3 = 3 18 != 3

OpenStudy (kittiwitti1):

Wait, what? @CoconutJJ

OpenStudy (coconutjj):

6/(1/3) = 6 X 3

OpenStudy (kittiwitti1):

...I got \[x=\frac{1}{2}\]I think it's wrong, let me try another approach. I did not simplify the fraction\[\frac{6}{\frac{1}{3}}\]

OpenStudy (coconutjj):

Yes x=1/2... 9x2 = 6x3

OpenStudy (kittiwitti1):

Oh. Okay. I was double-checking. So @Lindsay_552, the answer you got was not right...

OpenStudy (coconutjj):

You correct, @kittiwitti1

OpenStudy (kittiwitti1):

Thank you @CoconutJJ

OpenStudy (anonymous):

oh okay lol confussing thanks both of you!

OpenStudy (coconutjj):

u welcome..

OpenStudy (kittiwitti1):

Are you still confused? :c

OpenStudy (kittiwitti1):

I can show you the process I went through if you need it. c:

OpenStudy (anonymous):

nope i got it thanks!

OpenStudy (kittiwitti1):

Alright ^-^

OpenStudy (kittiwitti1):

You're very welcome~